what volumes of 0.500 M HNO (Ka = 4:0x104) and 0.500 M NANO must be mixed to Prepare lool of a solution buffered pH 3.55?
Answer #1

equation we from Henderson - Hasselbalch know , pH = pka + log [salt [acid] [NaNO2] O, 355 = -log(4.0x1074) + 20g HNOL} = 3.4